OpenStudy (anonymous):

John wants to rent a tent to go camping. The cost to rent a tent from Woodland Outfitters is a flat rate of $10.00 plus $0.50 for each night the tent is rented. Part a. Let x represent the number of nights John keeps the tent and write a cost function, c(x), for the cost to rent a tent from Woodland Outfitters. Part b. How much will John owe if he uses the tent for five nights?

OpenStudy (nikato):

How would you write the expression for "$0.50 for each night if x represents each night"

OpenStudy (ccswims):

A: c(x) = 0.50x + 10 B: c(x) = 0.50 * 5 +10 c(x) = 25 + 10 c(x) = $35